Pirate Jack is in a speed boat cruising along. He is going to jump off the boat at 7 m/s and onto the tube at the same height as the boat and being pulled 4 m behind. What angle does he have to jump at?
1
Expert's answer
2019-10-18T10:06:10-0400

The range of projectile:


"d=\frac{v^2}{g}\sin{2\theta}"

"4=\frac{7^2}{9.8}\sin{2\theta}\to\sin{2\theta}=0.8"

"\theta =0.5\arcsin{0.8}=27\degree"